Circular No. 5842

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J0422+32 A. V. Filippenko and T. Matheson, University of California, Berkeley,

report that the brightness of GRO J0422+32 appears to have increased significantly since the last reports in April (cf. IAUC 5761, 5764). Preliminary calibration of spectra (range 310-1000 nm) obtained on Aug. 13 UT with the Shane 3-m reflector at Lick Observatory yields V = 14.7 ± 0.5. Strong, broad Balmer absorption lines are visible in the rather blue continuum; Hα exhibits an emission component, as well. The object may have experienced a new accretion event at some time during the past four months.

A. J. Castro-Tirado, Danish Space Research Institute, Lyngby; and J. L. Ortiz, Instituto de Astrofisica de Andalucia, Granada, report: “We have been observing the optical counterpart of the soft x-ray transient GRO J0422+32, discovered in 1992 Aug. (cf. IAUC 5580, 5588, 5597). CCD frames taken at the Spanish 1.52-m telescope at the German-Spanish Calar Alto Observatory have shown that GRO J0422+32 has undergone a dramatic and unexpected increase in brightness by 3.6 mag in the last three days. Preliminary magnitudes (± 0.2): Aug. 10.105 UT, V = 18.7; 13.17, 15.1. The object may be emitting strongly at high energies, therefore xrays and γ-ray observations should be very valuable. Spectroscopic and photometric observations at all wavelengths are encouraged.”

SUPERNOVA 1993R IN NGC 7742 Filippenko and Matheson also report: “Careful inspection of fully cali-

brated spectra (range 310-1020 nm) obtained on June 28, July 14, and July 28 UT with the Shane 3-m reflector suggests that SN 1993R (cf. IAUC 5812) is a peculiar type-I supernova, rather than type II as stated on IAUC 5818. The line previously attributed to Hα is seen in type-Ia supernovae, and is generally thought to be [Fe II]. However, in other respects there are substantial differences from the late-time (nebular) spectra of normal type-Ia supernovae. The spectrum of SN 1993R bears some resemblance to that of the peculiar, subluminous type-Ia SN 1991bg, but the Ca II near-infrared triplet is strong in the former and very weak in the latter. The main difference from the nebular spectra of type-Ib and -Ic supernovae is that the [O I] 630-nm emission is weak in SN 1993R. The absolute R magnitude of SN 1993R at the time of discovery (probably near maximum brightness) was only –16 ± 1 for H_{0} = 50 km/s/Mpc. Note that the object is superposed on a very bright H II region.”

1993 August 13 Daniel W. E. Green